# Signature of preformed pairs in angle-resolved photoemission spectroscopy

### v1

@@ Abstract
We use density matrix renormalization group (DMRG) and variational exact diagonalization (VED) to calculate the single-electron removal spectral weight for the Hubbard-Holstein model at low electron densities. Tuning the strength of the electron-phonon coupling and of the Hubbard repulsion allows us to contrast the results for a liquid of polarons versus a liquid of bipolarons. The former shows spectral weight up to the Fermi energy, as expected for a (uncorrelated) metal. The latter has a gap in its spectral weight, set by the bipolaron binding energy, although this is also a (strongly correlated) metal. This difference suggests that angle-resolved photoemission spectroscopy could be used to identify liquids of pre-formed pairs. Furthermore, we show that the liquid of bipolarons is well approximated by an ensemble of bosons that are hard-core in momentum space, filling the states inside the Fermi sea but otherwise non-interacting. This new proposal for a strongly-correlated many-body wavefunction opens the way for studying various other properties of liquids of pre-formed pairs.
@@
Direct experimental detection of such preformed pairs remains a challenge. The extraordinary success of angle-resolved photoemission spectroscopy (ARPES) for characterizing single quasiparticle properties, has inspired a considerable effort to find generalizations that might provide direct spectroscopic signatures of pair formation. So far, the focus has been on detecting a pair of ejected electrons, whether from one 2e1 ; 2e1a ; 2e1b ; 2e1c or two 2e2 ; 2e2a incident photons. While these ideas are promising, such higher order processes are more difficult to detect.
In this Letter, we show that ARPES already contains clear signatures of preformed pair formation. We prove this by studying a one-dimensional (1D) Hubbard-Holstein model using density matrix renormalization group (DMRG)Schollwock2011 . Changing the ratio of the electron-phonon coupling vs. the electron-electron Hubbard repulsion allows us to vary the ground-state from a liquid of unbound polarons (quasiparticles whereby the carrier is dressed by a cloud of phonons) to a liquid of bipolarons (bound pairs of polarons). The DMRG calculated spectral weights prove that the latter is signalled by the appearance of a ’gap’ (set by the binding energy) below the Fermi energy, even though this phase is metallic. We explain this behavior using a simple theoretical picture of the bipolaron liquid as a liquid of hard-core bosons in momentum space, which is validated by comparison against the variationally exact DMRG results.
We emphasize that we expect these results to be generic. Using the Holstein electron-phonon coupling is convenient because of the multiple tools already developed to study it Wellein1997 ; Jeckelmann1998 ; Bonca1999 ; Barisic2002 ; Hohenadler2003 ; Berciu2006 ; Mitric2022 ; Mitric2023 , but the phenomenology remains similar for other types of bosons/couplings. Similarly, we do not expect dimensionality to play a role in this phenomenology; we consider a 1D chain simply for access to easier-to-converge DMRG calculations.
Figure 1: Spectral weight $A(k,\omega)$ for a system with $N_{e}=2$ electrons, at fixed $t=1$, $\Omega=2,\lambda=1$ and increasing $U=0,\dots,4$ (left to right panels). The white horizontal dashed line marks the chemical potential $\mu=E_{P}(0)$, while the black dashed line shows the inverted polaron dispersion, $-E_{P}(k)$ (up to an overall shift). As $U$ increases, the $N_{e}=2$ GS evolves from a bipolaron (spectra display a ’gap’ between the lowest binding energy feature and $\mu$) to two unbound polarons (spectral weight starting at $\mu$). See text for more details.
@@
where $c^{\dagger}_{i\sigma}$ and $b_{i}^{\dagger}$ create an electron with spin $\sigma$ and a boson, respectively, at site $i=1,\dots,N$; we set the lattice constant $a=1$. The number operator for electrons with spin-$\sigma$ is ${\hat{n}}_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}$, and ${\hat{n}}_{i}=\sum_{\sigma}{\hat{n}}_{i\sigma}$. The first term describes the bare electron hopping, the second is the on-site Hubbard repulsion, the third is the energy of the Einstein bosons ($\hbar=1$), and the last is the Holstein coupling, hereafter characterized by the effective electron-phonon coupling $\lambda=g^{2}/(2t\Omega)$.
We study this model (i) for systems with $N_{e}=1,2$ electrons on a finite chain with periodic boundary conditions employing variational exact diagonalization (VED) Bonca1999 ; Bonca2000 ; Ku2002 (we verified that the results are converged and correspond to the thermodynamic limit $N\rightarrow\infty$), and (ii) for finite electron concentrations $N_{e}/N=n_{\rm dop}<1$ in finite-size chains with open boundaries, using DMRG directly in the frequency space, with a recently developed variant Nocera2022 of the correction-vector approach Kuhner1999 ; Jeckelmann2002 ; NoceraPRE2016 . As customary AndreaRMP , we assume the ARPES intensity to be proportional to the spectral weight:
@@
Results: We first consider an infinite chain with $N_{e}=2$ electrons in its GS, to see the change in $A(k,\omega)$ when the two carriers are bound (at large $\lambda$, small $U$) vs. unbound (at small $\lambda$, large $U$). Either way, upon photo-ejection of an electron with momentum $k$, the system is left with one electron and a total momentum $-k$. The spectrum $E_{\alpha}^{(1)}$ is well known, see Fig. S1 in the Suppl. Mat. SM : the lowest-energy feature is the polaron band with energy $E_{P}(k)$, located below the polaron+one-phonon continuum starting at $E_{P}(0)+\Omega$, and other, even higher energy features Bonca1999 ; Berciu06 ; Bonca2019 . From Eq. (2), we therefore expect the feature with the lowest binding energy in $A(k,\omega)$ to appear at $\omega(k)=E_{GS}^{(2)}-E_{P}(k)\leq\mu$, and thus to disperse like the inverted polaron band. This is true independent of the nature of $|GS,2\rangle$; the latter is relevant only in setting the energy position of this feature.
Specifically, if the phonon mediated attraction is weak compared to $U$, the $N_{e}=2$ GS has two unbound polarons and $E_{GS}^{(2)}=2E_{P}(0)$. Otherwise the two electrons bind into a bipolaron and $E_{GS}^{(2)}=2E_{P}(0)-\Delta$, where $\Delta$ is the bipolaron binding energy Bonca2000 . It follows that one can infer whether the two carriers were bound from the location of the lowest binding energy feature in the ARPES spectral weight. This is confirmed in Fig. 1, where we plot $A(k,\omega)$ at fixed $\lambda=1,\Omega=2t$ for increasing values $U/t=0,\dots,4$. The dashed horizontal line marks $E_{P}(0)$ extracted from single-polaron calculations, which equals the chemical potential $\mu$ in this canonical calculation. Note that if $A(k,\omega)$ of Eq. (2) was evaluated in the grand-canonical ensemble ${\cal H}\rightarrow{\cal H}-\mu N_{e}$, this horizontal dashed line would shift to the customary ’Fermi energy’ $\omega=0$; we perform this shift for the latter figures.
@@
Electron-removal weight is indeed seen only below this line. (The vertical ’flashlight’ visible above it is due to a finite $\eta$ in $A(k,\omega)$). In all cases, the dispersion of the lowest binding energy feature indeed follows $-E_{P}(k)$ (up to a vertical shift). The latter is shown as the black dashed line extracted from the single polaron band, see Fig. S1 in SM . For $U\leq 2$, this lowest binding energy feature lies at a finite energy below the dashed $\omega=\mu$ line, as if the system is gapped (“insulating”). As discussed above, this gap equals the bipolaron binding energy, as we confirmed by separately computing $\Delta$. For $U\gtrsim 3$ the Hubbard repulsion is too strong and the ground-state is instead a “metal” consisting of two free polarons (we use quotation marks because the electron density is zero). Here, the ARPES spectral weight starts at the Fermi energy, as expected in a normal metal.
@@
Figure 2: The upper three rows show $A(k,\omega)$ calculated with DMRG for $n_{\rm dop}=0.0833$ and $t=1,\Omega=2$. The onsite repulsion $U=0\rightarrow 4$ increases from left to right, while the effective el-ph coupling $\lambda=0.25,0.5,1$ increases from top to bottom. The lowest row shows ARPES data assembled from single pair results like those of Fig. 1, as explained in the text. In all panels, the dashed horizontal line is the Fermi energy $\omega=0$, while vertical dashed lines indicate the Fermi momenta ±$\pm k_{F}$.
@@
Figure 1 proves that in $A(k,\omega)$, the key difference between having (or not) a pre-formed pair in the GS is the appearance (or not) of a gap below the Fermi energy. This is the first main result of this work. Before continuing, we comment on a secondary but important difference between the two cases. The left (bipolaron) panels show spectral weight tracing the inverted polaron dispersion at all $k$, whereas in the right (unbound polarons) panels, the spectral weight is strongly concentrated at $k=0$. This is expected. The bipolaron GS wavefunction $|BP\rangle\sim[\sum_{k}\phi_{k}c^{\dagger}_{k\uparrow}c^{\dagger}_{-k\downarrow}+\sum_{k,q}\phi_{k,q}c^{\dagger}_{k\uparrow}c^{\dagger}_{-k-q,\downarrow}b^{\dagger}_{q}+\dots]|0\rangle$ describes binding in real space. As a result, $c_{k\sigma}|BP\rangle\sim[\phi_{k}c^{\dagger}_{-k,-\sigma}+\sum_{q}\phi_{k,q}c^{\dagger}_{-k-q,-\sigma}b^{\dagger}_{q}+\dots]|0\rangle$ has a finite overlap with the single polaron eigenstate $|P,-k,-\sigma\rangle=[\sqrt{Z_{k}}c^{\dagger}_{-k,-\sigma}+\sum_{q}\psi_{k,q}c^{\dagger}_{-k-q,-\sigma}b^{\dagger}_{q}+\dots]|0\rangle$ for all $k$ values ($Z_{k}$ is the quasiparticle weight). In contrast, the unbound polarons GS is described well by the product of spin-up and spin-down polarons $|P,0,\uparrow\rangle\otimes|P,0,\downarrow\rangle$. The ARPES weight now is like that from of a single polaron Bonca2019 ; Bonca2022 with a strong peak at $\omega=\mu,k=0$, plus spectral intensity at $\omega=\mu-n\Omega$ for all $k$, indicating the $n\geq 1$ phonons left behind after the electron was photoejected. Indeed, we verified that the ARPES weight for the $N_{e}=1$ polaron GS is identical to that in the right panel of Fig. 1, proving that the second polaron is just a bystander.
@@
Next, we demonstrate that similar ARPES spectra are obtained for low, finite electron concentrations $n_{\rm dop}$. The top three rows of Figure 2 show ARPES spectra calculated with DMRG at $n_{\rm dop}=0.0833$, for $\lambda=0.25,0.5,1$ (top to bottom, respectively). Columns show $U/t=0\rightarrow 4$ (left $\rightarrow$ right, respectively). In all cases the spectral weight of the lowest binding-energy feature again follows $-E_{P}(k)$, see dashed black lines in the first column. A gap equal to the bipolaron binding energy is observed in the panels with small $U$ and large $\lambda$, where the GS is a liquid of bipolarons. We emphasize that this gap does not imply an insulating ground state: the bipolaron liquid is a metal, but a strongly correlated one. The gap closes as $U$ increases and/or $\lambda$ decreases and the bipolarons unbind, forming a liquid of polarons with momenta $k\in[-k_{F},k_{F}]$. Indeed, the ARPES weight for the lowest binding energy feature vanishes for $|k|>k_{F}$, as expected for a normal metal. Higher binding energy features again indicate the multi-phonon sidebands at $\omega=-n\Omega$.
@@
Figure 3: Comparison between ARPES spectra calculate with DMRG for $x=0.166,0.250,0.333$ (first row, left to right) and those generated from single pair calculations for a liquid of hard-core bosons (second row). Other parameters are $U=0,t=1,\Omega=2,\lambda=1$. See text for more details.
@@
The third row panels look similar to those shown for the same parameters but $n_{\rm dop}=0$ in Fig. 1. The key difference is that at $n_{\rm dop}>0$, the lowest binding energy feature is flat for all $|k|\leq k_{F}=\pi n_{\rm dop}/2$ (marked by vertical dashed lines), see also Fig. 3. This demonstrates that these bipolarons do not all occupy the bipolaron GS with total momentum $K=0$, as would be expected for weakly-interacting bosons. Instead, we find that we can mimic $A(k,\omega)$ for a finite concentration using the ARPES spectra calculated like in Fig. 1 for a single pair of electrons, if we assume that the bipolarons are hard-core bosons in momentum space, i.e. each momentum $|K|\leq k_{F}$ is occupied by one bipolaron, enforcing the Pauli exclusion principle for their fermions. Apart from this, the bipolarons are weakly interacting, given their low density. ARPES ejection of an electron with momentum $k$ can then be considered as the sum of ARPES spectral weight for ejecting an electron with momentum $k$ from a single bipolaron, summed over the possible bipolaron momenta $|K|\leq k_{F}$. Such calculations are shown in the lowest row of Fig. 2, and are in excellent quantitative agreement with the DMRG results for the same parameters, shown in the row just above it.
@@
This agreement persists up to much higher electron concentrations, as demonstrated in Fig. 3 by comparing the DMRG spectra obtained for $n_{\rm dop}=0.166,0.250$ and $0.333$ with spectra generated from sums of single pair VED results, as described above. The two sets of spectral weights are very similar. Both show the flattening of the lowest binding energy feature for $|k|\leq k_{F}$, as well as an overall broadening of this entire feature with increasing $n_{\rm dop}$. The VED results explain that this is a consequence of the fact that removing an electron with momentum $k$ from a bipolaron with momentum $K$ leaves behind an electron with momentum $K-k$, whose energy (and therefore binding energy) varies when summing up overall bipolaron contributions, $|K|\leq k_{F}$. The DMRG results also show this broadening but on a smaller scale, suggesting that the VED approximation misses some additional renormalization with increasing $n_{\rm dop}$ values.
@@
This demonstration that the liquid of bipolarons is well approximated by an ensemble of hard-core (in momentum-space) bosons with momenta $|K|\leq k_{F}$ but otherwise very weakly interacting, is the second major result of our work. It provides a simple physical picture for understanding this correlated unusual metal of pre-formed pairs. This allows us to speculate, for instance, on how this gap closes at a temperature comparable with the bipolaron binding energy. Above this temperature all pairs are unbound, hence we expect a typical metallic ARPES spectral weight starting at the Fermi energy (no gap). As the temperature is lowered and an increasingly larger fraction of polarons bind into pairs, their contribution to the spectral weight becomes gapped. As a result, the total spectral weight should be consistent with the opening of a pseudogap for $k_{B}T<\Delta$.
@@
Support for the speculations in the paragraph above is provided in Fig. S2 SM , although more work is needed to validate it. We note that pseudogaps are indeed observed in phases hypothesized to be liquids of pre-formed pairs, e.g. in underdoped cuprates. We emphasize that we do not claim that phonons are responsible for cuprate superconductivity. Instead, our claim is that any liquid of preformed pairs would exhibit a pseudogap in its ARPES weight at finite temperature, independent of the specific type of boson responsible for binding the carriers.
@@
Experimentally, this could be tested (i) with time resolved ARPES in materials with (pseudo)gaps attributed to pre-formed pairs AD2 . If pumping breaks some pairs, the time-scale over which the (pseudo)gap reopens plus the direct measurement of $\Delta$, might suffice to both verify the existence of pre-formed pairs, and to identify the boson(s) involved in the pairing; (ii) Applying hydrostatic pressure and/or uniaxial stress can decrease the $U/t$ ratio nn . In a material with sufficiently large $\lambda$ this could induce a crossover from a normal metal to a liquid of pre-formed pairs that could be observed in ARPES; (iii) trARPES could also be combined with experiments that endeavor to change the state of a system through optical pumping of specific phonon modes. Appearance of superconductivity well above the equilibrium $T_{C}$ has been reported in a variety of such systems pp ; pp1 ; pp2 ; pp3 ; pp4 . The opening of a (pseudo)gap in trARPES spectra would verify that preformed pairs are forming, an important first step towards confirming superconductivity. Finally, (iv) the evolution of ARPES spectra could most directly be verified in quantum simulators of Hubbard-Holstein models HHqs ; HHqsa , where parameters can be individually tuned.
@@
The above suggestions assume that the opening of a (pseudo)gap would be visible in trARPES data, which (i) might be challenging, and (ii) even if observed, would not rule out other possible explanations. The ansatz we identified here, that the liquid of pre-formed pairs is described well by hard-core bosons in momentum space, opens the way to calculate, in future work, its response to other probes, potentially resulting in the discovery of other fingerprints of this type of correlated state.

### v3

@@ Abstract
We use density matrix renormalization group (DMRG) and variational exact diagonalization (VED) to calculate the single-electron removal spectral weight for the Hubbard-Holstein model at low electron densities. Tuning the strength of the electron-phonon coupling and of the Hubbard repulsion allows us to contrast the results for a liquid of polarons versus a liquid of bipolarons. The former shows spectral weight up to the Fermi energy, as expected for a metal. The latter has a gap in its spectral weight, set by the bipolaron binding energy, although this is also a (strongly correlated) metal. This difference suggests that angle-resolved photoemission spectroscopy could be used to identify liquids of pre-formed pairs. Furthermore, we show that the one-dimensional liquid of incoherent bipolarons is well approximated by a ’Bose sea’ of bosons that are hard-core in momentum space, occupying the momenta inside the Fermi sea but otherwise non-interacting. This new proposal for a strongly-correlated many-body wavefunction opens the way for studying various other properties of incoherent (non-superconducting) liquids of pre-formed pairs in any dimension.
@@
Direct experimental detection of such preformed pairs remains a challenge, as is the formulation of a theoretical description of such a state. The extraordinary success of angle-resolved photoemission spectroscopy (ARPES) for characterizing single quasiparticle properties, has inspired a considerable effort to find generalizations that might provide direct spectroscopic signatures of pair formation. So far, the focus has been on detecting a pair of ejected electrons, whether from one 2e1 ; 2e1a ; 2e1b ; 2e1c or two 2e2 ; 2e2a incident photons. While these ideas are promising, such higher order processes are more difficult to detect.
In this Letter, we show that ARPES already contains clear signatures of preformed pair formation. We prove this by studying a one-dimensional (1D) Hubbard-Holstein model using density matrix renormalization group (DMRG)Schollwock2011 . Changing the ratio of the electron-phonon coupling vs. the electron-electron Hubbard repulsion allows us to vary the ground-state from a liquid of unbound polarons (quasiparticles whereby the carrier is dressed by a cloud of phonons) to a liquid of bipolarons (bound pairs of polarons). The DMRG calculated spectral weights prove that the latter is signalled by the appearance of a ’gap’ (equal to half the binding energy) below the Fermi energy, even though this phase is metallic. We explain this behavior using a simple theoretical picture of the bipolaron liquid as a ’Bose sea’ of hard-core bosons in momentum space, which is validated by comparison against the variationally exact DMRG results.
We emphasize that we expect this ARPES signature of pre-formed pairs to be generic. The Holstein coupling is convenient because of the multiple tools already developed to study it Wellein1997 ; Jeckelmann1998 ; Bonca1999 ; Barisic2002 ; Hohenadler2003 ; Berciu2006 ; Mitric2022 ; Mitric2023 , but we expect the polaron and bipolarons phenomenology to remain similar for other types of bosons and electron-boson couplings in all dimensions. Similarly, we do not expect dimensionality to play a role in the proposed ansatz for the liquids of incoherent pre-formed pairs, except that in higher-D these would occur at finite temperatures. Here we consider a 1D chain for access to easier-to-converge DMRG calculations and because its ground-state cannot be ordered (superconducting) Hoh ; Mermin-Wagner .
Figure 1: Spectral weight $A(k,\omega)$ for a system with $N_{e}=2$ electrons on $N=29$ sites, at fixed $t=1$, $\Omega=2,\lambda=1$ and increasing $U=0,\dots,4$ (left to right panels) computed with VED. The white horizontal dashed line marks the Fermi energy $\omega=\mu=E_{P}(0)-\Delta/2$ (see text for further discussion) while the black dashed line shows the inverted polaron dispersion, $-E_{P}(k)$ (up to an overall shift). As $U$ increases, the $N_{e}=2$ GS evolves from a bipolaron (spectra display a ’gap’ between the lowest binding energy feature and $\mu$) to two unbound polarons (spectral weight starting at $\mu$). See text for more details.
@@
where $c^{\dagger}_{i\sigma}$ and $b_{i}^{\dagger}$ create an electron with spin $\sigma$ and a boson, respectively, at site $i=1,\dots,N$, where $N$ is the number of sites and we set $a=1$, $\hbar=1$. The number operator for electrons with spin-$\sigma$ is ${\hat{n}}_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}$, and ${\hat{n}}_{i}=\sum_{\sigma}{\hat{n}}_{i\sigma}$. The first term describes the bare electron hopping, the second is the on-site Hubbard repulsion, the third is the energy of the Einstein phonons, and the last is the Holstein coupling, which we characterize in terms of the dimensionless electron-phonon coupling parameter $\lambda=g^{2}/(2t\Omega)$, equal to the ratio between the lattice deformation energy $g^{2}/\Omega$ in the atomic limit, and half the non-interacting bandwidth $2t$.
We study this model (i) for systems with $N_{e}=1,2$ electrons on a finite chain with periodic boundary conditions employing variational exact diagonalization (VED) Bonca1999 ; Bonca2000 ; Ku2002 (we verified that the results are converged and correspond to the thermodynamic limit $N\rightarrow\infty$), and (ii) for finite electron concentrations $N_{e}/N=n_{\rm dop}<1$ in finite-size chains with open boundaries, using DMRG directly in the frequency space, with a recently developed variant Nocera2022 of the correction-vector approach Kuhner1999 ; Jeckelmann2002 ; NoceraPRE2016 . As customary AndreaRMP , the ARPES intensity is assumed to be proportional to the spectral weight:
@@
Results: We first consider an infinite chain with $N_{e}=2$ electrons in its GS, to see the change in $A(k,\omega)$ when the two carriers are bound (at large $\lambda$, small $U$) vs. unbound (at small $\lambda$, large $U$). Either way, upon photo-ejection of an electron with momentum $k$, the system is left with one electron and total momentum $-k$. The spectrum $E_{\alpha}^{(1)}$ is well known, see Fig. 4 of End Matter: the lowest eigenstate is the polaron band with energy $E_{P}(k)$, located below the polaron+one-phonon continuum starting at $E_{P}(0)+\Omega$, and other, even higher energy features Bonca1999 ; Berciu06 ; Bonca2019 .
Using Eq. (2) and the fact that $E_{\alpha}^{(1)}\geq E_{P}(k)$, we expect the feature in $A(k,\omega)$ with the lowest binding energy to appear at energy $\omega(k)=E_{GS}^{(2)}-E_{P}(k)$, and thus it must disperse like the inverted polaron band $-E_{P}(k)$, regardless of whether the GS has two unbound polarons and $E_{GS}^{(2)}=2E_{P}(0)$, or a bipolaron and $E_{GS}^{(2)}=2E_{P}(0)-\Delta$, with $\Delta$ being the binding energy.
Furthermore, in this canonical calculation the chemical potential is $\mu=[E_{GS}^{(N_{e}+1)}-E_{GS}^{(N_{e}-1)}]/2$. If the GS has unbound polarons, then $\mu=E_{P}(0)$ and we find $\omega(k)\leq\omega(0)=\mu$, i.e. no gap is expected in the spectral weight. On the other hand, if the GS has a bipolaron then $\mu=E_{P}(0)-\Delta/2$ and we find $\omega(k)\leq\omega(0)=\mu-\Delta/2$, i.e. a gap equal to $\Delta/2$ is expected in the spectral weight.
These expectations are confirmed in Fig. 1, where we plot $A(k,\omega)$ at $\lambda=1,\Omega=2t$ for $U/t=0,\dots,4$. We shift energies so that the ’Fermi energy’ is at the usual ARPES $\omega-\mu$=0 value, shown by the horizontal line.
@@
Electron-removal weight is indeed seen only below $\omega-\mu$=0. (The vertical ’flashlight’ visible above is due to the finite $\eta$ in $A(k,\omega)$). In all cases, the dispersion of the lowest binding energy feature indeed follows $-E_{P}(k)$ (up to a vertical shift). The latter is shown by the black dashed line, and is extracted from the polaron band in Fig. 4. For $U\leq 2$, this lowest binding energy feature lies at a finite energy below the $\omega-\mu$=0 line, as if the system is gapped (“insulating”), and this gap indeed equals $\Delta/2$ as expected for a bipolaron GS. For $U\gtrsim 3$ the Hubbard repulsion is too strong and the ground-state is a “metal” of two unbound polarons (we use quotation marks because the electron density vanishes in the thermodynamic limit). The ARPES spectral weight starts at $\omega-\mu=0$, as expected for a ’metallic’ GS.
@@
Figure 2: The upper three rows show $A(k,\omega)$ calculated with DMRG for $n_{\rm dop}=0.0833$ and $t=1,\Omega=2$, $\eta=0.15$. DMRG simulations were performed on a $N=48$ sites chain with open boundary conditions and $N_{e}=4$. The onsite repulsion $U=0\rightarrow 4$ increases from left to right, while the effective el-ph coupling $\lambda=0.25,0.5,1$ increases from top to bottom. The lowest row shows $A(k,\omega)$ assembled from single pair VED results like those of Fig. 1, as explained in the text. Each single pair VED calculation was conducted on a system with $N=29$. For the VED simulations, $\eta$ was set to 0.05, while all other parameters remained consistent with those used in the DMRG calculations. In all panels, the dashed horizontal line is the Fermi energy $\omega=\mu$, while vertical dashed lines indicate the Fermi momenta ±$\pm k_{F}$.
@@
Figure 1 confirms that in $A(k,\omega)$, the key signature indicating the existence of a pre-formed pair (the bipolaron) is the appearance of a gap below the Fermi energy; this gap closes for unbound quasiparticles (polarons). This is the first main result of this work. Before continuing, we comment on a secondary but important difference between the two types of GS. The left (bipolaron) panels show spectral weight at all $k$ in the lowest-binding energy feature that traces (the shifted) $-E_{P}(k)$, whereas in the right (unbound polarons) panels, the spectral weight in this feature is strongly concentrated at $k=0$. This is because the bipolaron GS wavefunction $|BP\rangle\sim[\sum_{k}\phi_{k}c^{\dagger}_{k\uparrow}c^{\dagger}_{-k\downarrow}+\sum_{k,q}\phi_{k,q}c^{\dagger}_{k\uparrow}c^{\dagger}_{-k-q,\downarrow}b^{\dagger}_{q}+\dots]|0\rangle$ describes binding in real space. As a result, $c_{k\sigma}|BP\rangle\sim[\phi_{k}c^{\dagger}_{-k,-\sigma}+\sum_{q}\phi_{k,q}c^{\dagger}_{-k-q,-\sigma}b^{\dagger}_{q}+\dots]|0\rangle$ has a finite overlap at all $k$ values with the single polaron eigenstate with momentum $-k$ and spin $-\sigma$, defined as $|P,-k,-\sigma\rangle=[\sqrt{Z_{k}}c^{\dagger}_{-k,-\sigma}+\sum_{q}\psi_{k,q}c^{\dagger}_{-k-q,-\sigma}b^{\dagger}_{q}+\dots]|0\rangle$ ($Z_{k}$ is the quasiparticle weight). In contrast, the GS with unbound polarons is described well as an uncorrelated pair of GS polarons with opposite spins and zero momentum $|P,0,\uparrow\rangle\otimes|P,0,\downarrow\rangle$. The ARPES weight now is like that from of a single polaron Bonca2019 ; Bonca2022 with a strong peak at $\omega-\mu=0,k=0$, plus spectral intensity at $\omega-\mu=-n\Omega$ for all $k$, indicating the $n\geq 1$ phonons left behind after the electron was photoejected. Indeed, we verified that the ARPES weight for the $N_{e}=1$ polaron GS is identical to that in the right panel of Fig. 1, proving that the second polaron is just a bystander.
@@
Next, we demonstrate that similar ARPES spectra are obtained for low, finite electron concentrations $n_{\rm dop}$. The top three rows of Figure 2 show ARPES spectra calculated with DMRG at $n_{\rm dop}=0.0833$, for $\lambda=0.25,0.5,1$ (top to bottom, respectively). Columns correspond to $U/t=0\rightarrow 4$ (left $\rightarrow$ right, respectively). In all cases the spectral weight of the lowest binding-energy feature again follows $-E_{P}(k)$, see dashed black lines in the first column. A gap equal to $\Delta/2$ is observed in the panels with small $U$ and large $\lambda$, where the GS is a liquid of bipolarons. We emphasize that this gap does not imply an insulating ground state: the bipolaron liquid is a metal, albeit a strongly correlated one. The gap means that single carrier fluctuations are gapped, however two-carrier (pair) fluctuations are not. We verify this by showing that the Drude conductivity is finite for both types of liquids, see Appendix B of End Matter.
The gap in the single-electron removal spectral weight closes as $U$ increases and/or $\lambda$ decreases and the bipolarons unbind, forming a liquid of polarons with momenta $|k|\leq k_{F}$. Indeed, here the ARPES weight for the lowest binding energy feature vanishes for $|k|>k_{F}$, as expected for a normal metal. Higher binding energy features again reveal the multi-phonon sidebands at $\omega-\mu=-n\Omega$.
@@
Figure 3: Comparison between ARPES spectra calculated with DMRG for $n_{\rm dop}=0.166,0.250,0.333$ (first row, left to right) with those generated from single-pair VED calculations for a liquid of hard-core bosons at the closest attainable doping levels (second row). DMRG simulations were performed on a $N=48$ sites chain with open boundary conditions, while VED calculations were performed on a $N=29$ sites chain with periodic boundary conditions. Other parameters are $U=0,t=1,\Omega=2,\lambda=1,\eta=0.15$ (DMRG), $\eta=0.05$ (VED). See text for more details.
@@
The third row panels look similar to those shown for the same parameters but $n_{\rm dop}\rightarrow 0$ in Fig. 1. The key difference is that when $n_{\rm dop}>0$, the lowest binding energy feature is flat for all $|k|\leq k_{F}=\pi n_{\rm dop}/2$ (marked by vertical dashed lines). This is even more apparent in Fig. 3, which shows results for larger $n_{\rm dop}$. This demonstrates that these bipolarons do not all occupy the bipolaron GS with total momentum $K=0$, consistent with the fact that a superconducting GS is forbidden in 1D systems. Mermin-Wagner If macroscopic occupation of a single level (boson condensation) is forbidden, then the GS must consist of pairs occupying the low-lying bipolaron states with a microscopic occupation number. We find that we can mimic $A(k,\omega)$ for a finite concentration using the ARPES spectra calculated like in Fig. 1 for a single pair of electrons, if we assume that the bipolarons are hard-core bosons in momentum space, i.e. each momentum $|K|\leq k_{F}$ is occupied by one bipolaron. ARPES ejection of an electron with momentum $k$ can then be considered as the sum of ARPES spectral weight for ejecting an electron with momentum $k$ from a single bipolaron, summed over the possible bipolaron momenta $|K|\leq k_{F}$. Calculations for such ’Bose seas’ are shown in the lowest row of Fig. 2, and are in excellent quantitative agreement with the DMRG results for the same parameters, shown in the row just above it.
@@
This agreement persists up to much higher electron concentrations, as demonstrated in Fig. 3 by comparing the DMRG spectra obtained for $n_{\rm dop}=0.166,0.250$ and $0.333$ with spectra generated from sums of single pair VED results, as described above. The two sets of spectral weights are very similar. Both show the flattening of the lowest binding energy feature for $|k|\leq k_{F}$, as well as an overall broadening of this entire feature with increasing $n_{\rm dop}$. The VED results explain that this broadening is a consequence of the fact that removing an electron with momentum $k$ from a bipolaron with momentum $K$ leaves behind an electron with momentum $K-k$, whose energy (and therefore binding energy) varies when summing up overall bipolaron contributions, $|K|\leq k_{F}$.
@@
This demonstration that the liquid of bipolarons is well approximated by a ’Bose sea’ of hard-core (in momentum-space) bosons with momenta $|K|\leq k_{F}$ but otherwise very weakly interacting, is the second major result of our work. It provides a simple physical picture for understanding this correlated unusual 1D metal of pre-formed pairs, and allows us to speculate, for instance, on how this single-particle gap closes at a temperature comparable with the bipolaron binding energy. Above this temperature all pairs are unbound, hence we expect a typical metallic ARPES spectral weight starting at the Fermi energy (no gap). As the temperature is lowered and an increasingly larger fraction of polarons bind into pairs, their contribution to the spectral weight becomes gapped. As a result, the total spectral weight should be consistent with the opening of a pseudogap for $k_{B}T<\Delta$.
@@
Support for the speculations in the paragraph above is provided in Fig. LABEL:figC1_in_the_End_Matter, although more work is needed to validate it. We note that pseudogaps are indeed observed in higher-D phases hypothesized to be liquids of pre-formed pairs, e.g. in underdoped cuprates. We emphasize that we do not claim that phonons are responsible for cuprate superconductivity; the low carrier limit of our model does not support antiferromagnetic order, and its 2D bipolarons have $s$, not $d$-wave symmetry. Instead, our claim is that a liquid of preformed pairs would likely exhibit a pseudogap in its ARPES weight at finite $T$, regardless of the specific type of boson responsible for binding the carriers.
@@
Experimentally, this could be tested (i) with time resolved ARPES (trARPES) in materials with (pseudo)gaps attributed to pre-formed pairs AD2 . If pumping breaks some pairs, the time-scale over which the (pseudo)gap reopens plus the direct measurement of $\Delta$, might suffice to both verify the existence of pre-formed pairs, and to identify the boson(s) involved in the pairing; (ii) Applying hydrostatic pressure and/or uniaxial stress can decrease the $U/t$ ratio nn . In a material with sufficiently large $\lambda$ this could induce a crossover from a normal metal to a liquid of pre-formed pairs that could be observed in ARPES; (iii) trARPES could be combined with experiments that endeavor to change the state of a system through optical pumping of specific phonon modes. Appearance of superconductivity well above the equilibrium $T_{C}$ has been reported in a variety of such systems pp ; pp1 ; pp2 ; pp3 ; pp4 . The opening of a (pseudo)gap in trARPES spectra would verify that preformed pairs are forming, an important first step towards confirming superconductivity. Finally, (iv) the evolution of ARPES spectra could most directly be verified in quantum simulators of Hubbard-Holstein models HHqs ; HHqsa , where parameters can be individually tuned.
@@
The above suggestions assume that the opening of a (pseudo)gap would be visible in trARPES data, which (i) might be challenging, and (ii) even if observed, would not rule out other possible explanations.
The ’Bose sea’ of pairs ansatz we identified in 1D proposes a new type of microscopic wavefunction for liquids of uncondensed pre-formed pairs in any dimension, as a possible intermediary phase between a superconductor with a macroscopic concentration of condensed pairs, and a normal phase where all pairs are unbound. Our work suggests that if the binding energy $\Delta$ is considerable, thermal fluctuations may first excite pairs from the condensate into finite $K$ pair states. The coherence is lost at $T_{C}$ where the pairs spread over the Bose sea, leading to an unusual metal of pre-formed pairs (disordered-phase superconductor). The breaking of the pairs into single quasiparticles would occur at much higher temperatures $k_{B}T\sim\Delta$, leading to a normal metal.
Our ’Bose sea’ ansatz for this unusual metal of pre-formed pairs opens the way to calculate, in future work, its response to other probes, potentially providing other experimental fingerprints of this type of correlated state.